OpenStudy (anonymous):

write an equation of the line through the point parallel to the given line and perpendicular to the given line. Point: (-3,2). Line: x+y=7

OpenStudy (anonymous):

@ParthKohli @mathteacher1729 @amistre64

OpenStudy (amistre64):

do parallel lines have the same slope?

OpenStudy (anonymous):

yes..i think

OpenStudy (amistre64):

and its asking for a perp line thru the point given ; so whats a property of perp lines? slopewise that is

OpenStudy (anonymous):

it is opposite..like if the original is 2/3, it would be i believe -3/2

OpenStudy (amistre64):

its seems to be asking for both a perp and a parallel

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

good, 2/3 would be -3/2

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

well, lets find the slopd of the given line

OpenStudy (anonymous):

kk

OpenStudy (amistre64):

when given the form: ax + by = c the slope can be determined as -a/b

OpenStudy (anonymous):

y=-x+7?

OpenStudy (anonymous):

oh ok, so slope of line is -1?

OpenStudy (amistre64):

-1 is good, and thats going to be our parallel line slope; what our perp slope then?

OpenStudy (anonymous):

1

OpenStudy (amistre64):

good, now use point slope forms to establish line equations with

OpenStudy (anonymous):

slope of the given line is -1 slope of required line will be 1 so point slope form y-y1=m(x-x1) y-2=1(x-(-3) x-y+5=0

OpenStudy (anonymous):

wait what

OpenStudy (amistre64):

point slope form: y-Py = m(x-Px)

OpenStudy (anonymous):

Can u explain? first the parallel ?

OpenStudy (anonymous):

the equation

OpenStudy (amistre64):

your given the point (Px,Py): (-3,2)

OpenStudy (amistre64):

your prolly used to seeing it as: y-y1 = m(x-x1)

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

plug in our values: parallel = same slope: y-2 = -1(x--3) perp slope = y-2 = 1(x--3)

OpenStudy (anonymous):

ok yea

OpenStudy (amistre64):

two negatives equals a positive ;)

OpenStudy (anonymous):

Ok so my answer: parallel: y-2=-1(x+3) and perp: y-2=1(x+3)

OpenStudy (anonymous):

complete solution is there in my response take a look.

OpenStudy (amistre64):

thats correct

OpenStudy (anonymous):

so thats my answ? :)

OpenStudy (anonymous):

x-y+5=0 is answer

OpenStudy (anonymous):

thx @sami-21 and @amistre64

OpenStudy (amistre64):

as long as there isnt some overarching paragraph the says: put this in "such and such" form; its fine